# WcDT: World-centric Diffusion Transformer for Traffic Scene Generation

### v3

@@ Abstract
In this paper, we introduce a novel approach for autonomous driving trajectory generation by harnessing the complementary strengths of diffusion probabilistic models (a.k.a., diffusion models) and transformers. Our proposed framework, termed the "World-Centric Diffusion Transformer"(WcDT), optimizes the entire trajectory generation process, from feature extraction to model inference. To enhance the scene diversity and stochasticity, the historical trajectory data is first preprocessed into "Agent Move Statement" and encoded into latent space using Denoising Diffusion Probabilistic Models (DDPM) enhanced with Diffusion with Transformer (DiT) blocks. Then, the latent features, historical trajectories, HD map features, and historical traffic signal information are fused with various transformer-based encoders that is used to enhance the interaction of agents with other elements in the traffic scene. The encoded traffic scenes are then decoded by a trajectory decoder to generate multimodal future trajectories. Comprehensive experimental results show that the proposed approach exhibits superior performance in generating both realistic and diverse trajectories, showing its potential for integration into automatic driving simulation systems. Our code is available at https://github.com/yangchen1997/WcDT.
Figure 1: World-centric model and agent-centric model: (a) The conventional "Agent-centric model" is common in trajectory prediction, including on the Sim Agents leaderboard. (b) Our approach replaces complex coordinate transformations with position embeddings, enhancing efficiency in multi-scenario, multi-agent trajectory generation.
@@ I Introduction
[…]
Autonomous driving is a transformative technology aimed at reducing driver fatigue and traffic congestion by enabling autonomous vehicle operation [1, 2, 3, 4]. Developing these algorithms involves iterative optimization for safety and performance [5, 6, 7], but real-world testing poses challenges due to time constraints, safety concerns, regulatory hurdles, and high costs [8, 9]. Simulators play a vital role in the cost-effective testing and evaluation of autonomous driving systems (ADS) by providing controllable environments [10, 11]. To be effective, they must realistically replicate traffic scenarios and driver behaviors [12]. Current simulators rely on replaying driving logs or heuristic controllers [13, 12], limiting diversity and unpredictability in real-world behavior, which impacts ADS validation [14]. Multimodal motion prediction approaches [15, 16, 17, 18, 19, 20] have shown promise in traffic scene generation but struggle to generate diverse actions for all agents using comprehensive global information [15]. Generative adversarial networks (GANs) and Variational Auto-Encoders (VAEs) have been applied to traffic scene generation but face limitations. These models often lack diversity, reflecting training data distributions [12], and GANs suffer from unstable adversarial training [14]. Additionally, they fail to capture agent trajectory smoothness, leading to unrealistic results [14], and typically focus on individual vehicle paths, neglecting all agents. Recently, diffusion models have emerged as a promising alternative for diverse traffic scenarios [21, 12, 14, 22], treating generation as an inverse diffusion process. However, these models often require agent-centric Cartesian coordinates [15] and generate only one trajectory per agent per inference. In this paper, we propose a novel framework for traffic scene generation tailored to autonomous driving, leveraging diffusion models and transformer-based encoder-decoder architectures. Our "World-Centric Diffusion Transformer" (WcDT) framework optimizes trajectory generation from feature extraction to inference, enabling coherent and joint future movements for various agents in a single inference. Our contributions include:
@@ III-A Traffic scene representation
  • $\mathcal{D}$: Dimensionality of feature representations.
@@
  • Agent move statement and features: To mitigate the impact of varying agent positions on historical and future trajectories, we introduce absolute states for past and prospective agent states. For agent $i$ at time step $t$, the state $s_{t}^{i}$ is defined as $s_{t}^{i}=[(x_{t}-x_{t-1}),(y_{t}-y_{t-1}),(\theta_{t}-\theta_{t-1}),(v_{t}-v_{t-1})]$, where $x_{t}$, $y_{t}$, $\theta_{t}$, and $v_{t}$ represent longitudinal position, lateral position, heading angle, and velocity, respectively. The feature space for each agent is $[\mathcal{A},\mathcal{T}_{h}-1,\mathcal{D}]$.
@@
  • Map feature: The map features, denoted as $[1,\mathcal{L},\mathcal{P},\mathcal{D}_{m}]$, encompass key lane details in a traffic scenario, including positions and types. Each lane line $l_{t}\in\mathcal{L}$ at the current time step is represented using one-hot encoding to specify the positions of all points along the lane and its type.
@@ III-B Action Diffusion
To enhance trajectory diversity in WcDT, we encode agent actions into latent space to increase variability. These latent features are then input into the scene encoder. We use Denoising Diffusion Probabilistic Models (DDPM) [30] for action encoding. Although DDPM traditionally employs U-Net architectures, recent research [31] demonstrates that transformers can achieve comparable performance without U-Net’s inductive biases. Consequently, we replace U-Net with Diffusion Transformers (DiTs) to improve performance and ensure diverse agent trajectories. Figure 3 shows the architecture of conditional DiT blocks for encoding "latent action features". The network takes random noise, time steps, and historical trajectories as inputs and produces latent action features for the scene encoder. The DDPM loss function guides the network to generate latent features consistent with agent kinematics [30], thus enhancing trajectory variability. The loss function for the DiT module is as follows:
@@ III-E Loss functions
III-F Loss function
The composite loss function combines multiple constraints and modalities, given by:
$\mathcal{L}_{total}=$ $\ \lambda_{diff}\cdot\mathcal{L}_{diff}+\lambda_{reg}\cdot\text{Huber}(\text{Traj}_{p},\text{Traj}_{gt})$
$\ +\lambda_{cls}\cdot\left(-\sum_{i=1}^{M}y_{i}\log(p_{i})\right)$ (13)
The loss structure, comprising diffusion loss $\mathcal{L}_{diff}$ (L2 norm between predicted and actual noise), regression loss $\mathcal{L}_{reg}$ (Huber loss between predicted and ground truth trajectories), and classification loss $\mathcal{L}_{cls}$ (improving modality selection accuracy), balanced by weighting factors $\lambda_{diff}$, $\lambda_{reg}$, and $\lambda_{cls}$, enables the model to effectively handle diverse agent behaviors for optimal trajectory predictions in dynamic traffic scenarios.
@@ IV-C Ablation studies on diffusion model
We evaluate the effect of the "latent action features" encoding, comparing random noise inputs, the Unet network, and our custom DiT block. As shown in Table II, the DiT module consistently achieves the lowest ADE and MinADE scores, along with the highest composite score. This demonstrates that the diffusion model with the DiT block significantly improves action diversity while maintaining realistic trajectories.

### v4

@@ Abstract
In this paper, we introduce a novel approach for autonomous driving trajectory generation by harnessing the complementary strengths of diffusion probabilistic models (a.k.a., diffusion models) and transformers. Our proposed framework, termed the “World-centric Diffusion Transformer”(WcDT), optimizes the entire trajectory generation process, from feature extraction to model inference. To enhance the scene diversity and stochasticity, the historical trajectory data is first preprocessed into “Agent Move Statement” and encoded into latent space using Denoising Diffusion Probabilistic Models (DDPM) enhanced with Diffusion with Transformer (DiT) blocks. Then, the latent features, historical trajectories, HD map features, and historical traffic signal information are fused with various transformer-based encoders that is used to enhance the interaction of agents with other elements in the traffic scene. The encoded traffic scenes are then decoded by a trajectory decoder to generate multimodal future trajectories. Comprehensive experimental results show that the proposed approach exhibits superior performance in generating both realistic and diverse trajectories, showing its potential for integration into automatic driving simulation systems. Our code is available at https://github.com/yangchen1997/WcDT.
Figure 1: World-centric model and agent-centric model: (a) The conventional ”Agent-centric model” is common in trajectory prediction, including on the Sim Agents leaderboard. (b) Our approach replaces complex coordinate transformations with position embeddings, enhancing efficiency in multi-scenario, multi-agent trajectory generation.
@@ I Introduction
Autonomous driving is a transformative technology aimed at reducing driver fatigue and traffic congestion by enabling autonomous vehicle operation [1, 2, 3, 4]. Developing these algorithms involves iterative optimization for safety and performance [5, 6, 7], but real-world testing poses challenges due to time constraints, safety concerns, regulatory hurdles, and high costs [8, 9]. Simulators play a vital role in the cost-effective testing and evaluation of autonomous driving systems (ADS) by providing controllable environments [10, 11]. To be effective, they must realistically replicate traffic scenarios and driver behaviors [12]. Current simulators rely on replaying driving logs or heuristic controllers [13, 12], limiting diversity and unpredictability in real-world behavior, which impacts ADS validation [14]. Multimodal motion prediction approaches [15, 16, 17, 18, 19, 20] have shown promise in traffic scene generation but struggle to generate diverse actions for all agents using comprehensive global information [15]. Generative adversarial networks (GANs) and Variational Auto-Encoders (VAEs) have been applied to traffic scene generation but face limitations. These models often lack diversity, reflecting training data distributions [12], and GANs suffer from unstable adversarial training [14]. Additionally, they fail to capture agent trajectory smoothness, leading to unrealistic results [14], and typically focus on individual vehicle paths, neglecting all agents. Recently, diffusion models have emerged as a promising alternative for diverse traffic scenarios [21, 12, 14, 22], treating generation as an inverse diffusion process. However, these models often require agent-centric Cartesian coordinates [15] and generate only one trajectory per agent per inference. In this paper, we propose a novel framework for traffic scene generation tailored to autonomous driving, leveraging diffusion models and transformer-based encoder-decoder architectures. Our ”World-centric Diffusion Transformer” (WcDT) framework optimizes trajectory generation from feature extraction to inference, enabling coherent and joint future movements for various agents in a single inference. Our contributions include:
@@ III-A Traffic scene representation
  • $\mathcal{D}$: Dimensionality of different traffic elements in a traffic scenario ($\mathcal{D}_{a}$ represents the features of an agent, $\mathcal{D}_{t}$ represents the features of a traffic light, and $\mathcal{D}_{m}$ represents the features of a map element).
@@
  • Agent move statement and features: To mitigate the impact of varying agent positions on historical and future trajectories, we introduce absolute states for past and prospective agent states. For agent $i$ at time step $t$, the state $s_{t}^{i}$ is defined as $s_{t}^{i}=[(x_{t}-x_{t-1}),(y_{t}-y_{t-1}),(\theta_{t}-\theta_{t-1}),(v_{t}-v_{t-1})]$, where $x_{t}$, $y_{t}$, $\theta_{t}$, and $v_{t}$ represent longitudinal position, lateral position, heading angle, and velocity, respectively. The feature space for each agent is $[\mathcal{A},\mathcal{T}_{h}-1,\mathcal{D}_{a}]$.
@@
  • Map feature: The map features, denoted as $[1,\mathcal{L},\mathcal{P},\mathcal{D}_{m}]$, encompass key lane details in a traffic scenario, including positions and types. Each lane line $l_{t}\in\mathcal{L}$ at the current time step is represented positions of all points along the lane and using one-hot encoding to specify the its type.
@@ III-B Action Diffusion
To enhance trajectory diversity in WcDT, we encode agent actions into latent space to increase variability. These latent features are then input into the scene encoder. We use Denoising Diffusion Probabilistic Models (DDPM) [30] for action encoding. Although DDPM traditionally employs U-Net architectures, recent research [31] demonstrates that transformers can achieve comparable performance without U-Net’s inductive biases. Consequently, we replace U-Net with Diffusion Transformers (DiTs) to improve performance and ensure diverse agent trajectories. Figure 3 shows the architecture of conditional DiT blocks for encoding ”latent action features”. The network takes random noise, time steps, and historical trajectories as inputs and produces latent action features for the scene encoder. The DDPM loss function guides the network to generate latent features consistent with agent kinematics [30], thus enhancing trajectory variability. The loss function for the DiT module is as follows:
@@ IV-C Ablation studies on diffusion model
We evaluate the effect of the ”latent action features” encoding, comparing random noise inputs, the Unet network, and our custom DiT block. As shown in Table II, the DiT module consistently achieves the lowest ADE and MinADE scores, along with the highest composite score. This demonstrates that the diffusion model with the DiT block significantly improves action diversity while maintaining realistic trajectories.